# Estimating the Lateral Motion States of an Underwater Robot by Propeller Wake Sensing Using an Artificial Lateral Line

### v1

@@ IV-D Experimental Results
Figure 6: The experimental results of motion state estimation comparing the proposed hybrid CNN-BiLSTM model and the commonly-used CNN model. The adopted performance evaluation metrics are the RMSE for the lateral displacement and the ACC for the traveling speed and direction of motion.
@@
Figure 5 shows the experimental results of the lateral motion state estimation of the leader propeller. Table I presents the statistical estimation results for each motion state over all the experimental tests in Case 1 and Case 2, respectively. Specifically, Figs 5(a) and 5(b) show the estimates and the actual values of the lateral displacement of the traveling leader propeller for Case 1 and Case 2, respectively. The red dashed line represents perfect estimation for reference; the blue circles represent the state estimates $\hat{x}$. The experimental results show that the estimated displacements are generally close to the actual value indicating a good regression performance. Furthermore, Table I reveals that a larger longitudinal displacement leads to a slightly increased root mean square error (RMSE) at 0.07 compared to that of 0.05 from the smaller longitudinal displacement. Meanwhile, the standard deviation of the estimation error from both cases are satisfactorily small at 0.0042 and 0.0012, respectively. Figures 5(c)-5(f) show the estimation results regarding the travelling speed ${v}$ and the direction of motion ${d}$ of the leader propeller represented by the confusion matrices that demonstrate the classification performance. The rows and the columns of the confusion matrix correspond to the actual and predicted classes, respectively. The diagonal and off-diagonal elements correspond to the true and false observation results, respectively. Figures 5(c) and 5(d) illustrate the estimation performance regarding the direction of motion ${d}$. $P$ and $N$ represent along the positive and negative $y$-axis directions, respectively. It is observed that the accuracy of the direction state classification is greater than 99.5% for the smaller longitudinal displacement $y=250$ mm while the accuracy is greater than 99% for the larger longtudinal displacement $y=300$ mm. The RMSE is less than 0.2% for both the cases. Specifically, Figures 5(e) and 5(f) illustrate the estimation performance regarding the traveling speed ${v}$ where 5 different lateral speeds of $0.4,0.5,0.6,0.7,0.8$ m/min are used corresponding to the first to the fifth columns of the confusion matrix labeled as $1$ to $5$. We observe that the accuracy of the lateral speed classification is consistently greater than 91% with a mean square error less than 1.5%.
@@
Case
RMSE of $\hat{x}$
[…]
Case 1 (with Optimization Parameters)
0.0498$\pm$ 0.0042
97.37$\pm$0.72%
99.75$\pm$0.11%
Case 1 (with Random Parameters)
0.0661$\pm$ 0.0120
93.47$\pm$7.32%
98.31$\pm$1.86%
Case 2 (with Optimization Parameters)
0.0712$\pm$ 0.0012
95.58$\pm$1.21%
99.37$\pm$0.21%
Case 2 (with Random Parameters)
0.0645$\pm$ 0.0159
89.16$\pm$24.23%
98.89$\pm$1.9%
The satisfactory experimental results validate the proposed motion state estimation algorithm, which further indicate that the propeller wake features extracted from the ALL distributed pressure measurements are sufficiently rich for estimating the motion states of a leader propeller-driven underwater robot in a leader-follower formation.
IV-E Ablation Study
Ablation experiments are conducted to verify the proposed hybrid CNN-BiLSTM design in feature extraction. Specifically, the BiLSTM network is removed with only the CNN model left as the comparison. The comparison experimental results are shown in Fig. 6 in terms of the three estimated states, i.e., the lateral displacement $x$, the lateral speed $v$, and the direction of motion $d$ of the leader propeller. The green bars represent the mean of the estimates using the proposed CNN-BiLSTM model while the purple bars represent the mean of the estimates using the CNN model. The error bars represent the standard deviations of the estimates for all the three motion states of interest over all the experimental tests in Case 1 and Case 2, respectively. From the comparison results, we observe a consistent performance improvement using CNN-BiLSTM compared to CNN for all the three motion states of interest and two longitudinal displacement cases. Specifically, Fig. 6(a) shows a roughly 30% improvement in the RMSE of the lateral displacement regression. Figure 6(b) demonstrates a significantly enhanced accuracy of over 50% improvement in estimating the travelling speed of the leader propeller. Similarly, a superior performance is observed with an accuracy improvement of over 10% in estimating the direction of motion as shown in Fig. 6(c). The ablation study validates the necessity and the effectiveness of the proposed CNN-BiLSTM model particularly over the traditionally used CNN network.
@@ V-A Length of Time Series on Estimation Performance
Figure 7 shows the experimental results of the influence of the time-series sequence length on the estimation performance represented by the corresponding evaluation metrics. Specifically, the classification accuracy (ACC) for the traveling speed $v$ and the motion of direction $d$ follows the axis scale on the left hand side, and the regression root mean square error (RMSE) for the lateral displacement $x$ follows the axis scale on the right hand side. We observe that as the sequence length increases, the RMSE of the lateral displacement decreases slightly indicating a gentle performance improvement. On the other hand, the classification accuracy of both the travelling speed and the motion of direction increases considerably when the sequence length increases. Overall, the longer length of the time-series is used in the estimation algorithm, the higher the motion estimation performance is achieved, which we conjecture comes from the extra flow dynamics information embedded in the added sampled data. However, longer sequence length leads to more time delay in the estimation process, thus possibly hindering real-time control tasks. Comprehensively considering the estimation performance and the real-time requirement of the ALL sensing system, the selection of the sequence length of the sensor measurement time series is expected to be resolved with respect to different applications. Particularly, $sl=64$ is adopted in this paper.
@@ V-B Optimized Task Weights on Estimation Performance
Figure 8: The comparative experimental results of the statistical fitness $g$ between using optimized and random task weights in Case 1 and Case, respectively.
The loss function $l$ in Eq. (6) designed for training the estimator network is a linear combination of the corresponding performance loss for the estimation tasks regarding the three motion states of interest with weight $\lambda_{i}$ as the design parameters. Formulated as a multi-objective optimization problem for the multi-state estimation, a biomimetic algorithm WOA is adopted to find the optimal task weight $\lambda_{i}$. This section conducts comparative experiments between using the optimized weights and using random weights, aiming to shed light on the influence of the weight optimization on the overall estimation performance. The comparative experimental results are shown in Fig. 8 and Table I. From Fig. 8, we observe that the fitness $g$ with optimized weights in Eq. (7) is significantly reduced by roughly 57% and 17% in Case 1 and Case 2, respectively. To further analyze the difference between optimized weights and randomly generated weights, comparative experiments are repeated 10 times for 10 selected trained estimation models. Table I shows the statistical results of the evaluation performances. The first and third rows are the estimation results using the optimized weights and the second and fourth rows use randomly generated weights. From Table I, we observe an estimation accuracy improvement by 4% for the speed $v$ and 1.4% for the direction of motion as well as an estimation accuracy improvement (RMSE) by 24.6% for the displacement $x$ in Case 1. We want to point out the estimation result on displacement $x$ in Case 2 decreases by 0.6% which comes from the highly-coupling phenomenon of the multi-objective optimization. The standard deviations of the estimation accuracy for all the three states are reduced by at least 65% with the optimized weights, indicating a more robust and consistent estimation performance over all the motion states and the testing conditions.
@@
V-C Comparison with Other Wake Sensing Methods
This section compares our proposed method with some of the existing wake sensing methods to highlight its unique contributions and advantages in the field. Table II shows the comparison results in terms of the generation method of the wake flow and whether multiple motion states are estimated simultaneously. The table refers to the selected literature by their organizations including the Massachusetts Institute of Technology (MIT)[14], Peking University (PKU)[37, 38, 39, 40], Harbin Engineering University (HEU)[41], National Taiwan University (NTU)[13], and the University of Maryland (UMD)[42]. In contrast to the existing methods that primarily focus on flow fields generated by periodic body undulation and tail flapping of bioinspired robotic fish, our study addresses the dynamic and complex wake flow generated by underwater robots propelled by high-speed rotating propellers. Meanwhile, our method provides the simultaneous estimation of multiple motion states via the multi-output network design that combines regression and classification operations. The existing studies demonstrate various important applications of ALL-based wake sensing such as hydrodynamic efficiency enhancement[14], robot formation control[39, 40, 41], and flow-relative localization[37, 38, 13, 42]. Consequently, our method holds great potential for expanding these applications to a wider range of uses in propeller-driven underwater robots.
@@ VI Conclusions
This paper investigated the challenging problem of estimating the relative motion states of a propeller-driven leader underwater robot in the leader-follower formation. Making a hypothesis that the ALL pressure measurements provide partial but sufficiently rich information of the propeller wake of a leader underwater robot, we proposed a novel motion state estimation method that assimilated distributed pressure measurements sampled from the dynamic and complex propeller wake. Specifically, to estimate the lateral displacement, speed, and direction of motion of the leader propeller, a hybrid network that combined 1DCNN and BiLSTM was designed to extract the spatiotemporal characteristics of the flow field, followed by a multi-output network for state estimation. An experimental testbed was constructed. Extensive experimental tests were conducted, the results of which validated the proposed estimation method. Furthermore, comparative studies were conducted, providing insights into the design of the multi-objective loss function for training the estimator network as well as the selection of the sequence length of the pressure measurement time series used in the estimation algorithm.
In future work, we will apply the proposed wake flow sensing approach to investigate three-dimensional motion state estimation of a propeller-driven leader underwater robot. We also plan to integrate the estimation algorithm with feedback control design and experimentally test leader-follower formation control of propeller-driven underwater robots.

### v2

@@ IV-D Experimental Results
Figure 6: The experimental results of motion state estimation comparing the proposed method with traditional frequency analysis-based (STFT) and learning-basded (CNN) methods. Evaluated by the estimation accuracy in terms of all the three motion states of interest, the proposed method consistently outperforms the conventional methods from the literature in addressing the unique propeller wake estimation problem.
@@
Figure 5 shows the experimental results of the lateral motion state estimation of the leader propeller. Table I presents the statistical estimation results for each motion state over all the experimental tests in Case 1 and Case 2, respectively. Specifically, Figs. 5(a) and 5(b) show the estimates and the actual values of the lateral displacement of the traveling leader propeller for Case 1 and Case 2, respectively. The red dashed line represents perfect estimation for reference; the blue circles represent the state estimates $\hat{x}$. The experimental results show that the estimated displacements are generally close to the actual value indicating a good regression performance. Furthermore, Table I reveals that a larger longitudinal displacement leads to a slightly increased root mean square error (RMSE) at 0.07 compared to that of 0.05 from the smaller longitudinal displacement. Meanwhile, the standard deviation of the estimation error from both cases are satisfactorily small at 0.0042 and 0.0012, respectively.
Figures 5(c)-5(f) show the estimation results regarding the traveling speed ${v}$ and the direction of motion ${d}$ of the leader propeller represented by the confusion matrices that demonstrate the classification performance. The rows and the columns of the confusion matrix correspond to the actual and predicted classes, respectively. The diagonal and off-diagonal elements correspond to the true and false observation results, respectively. Figs. 5(c) and 5(d) illustrate the estimation performance regarding the direction of motion ${d}$. $P$ and $N$ represent along the positive and negative $y$-axis directions, respectively. It is observed that the accuracy of the direction state classification is greater than 99.5% for the smaller longitudinal displacement $y=250$ mm while the accuracy is greater than 99% for the larger longtudinal displacement $y=300$ mm. The RMSE is less than 0.2% for both the cases. Specifically, Figs. 5(e) and 5(f) illustrate the estimation performance regarding the traveling speed ${v}$ where 5 different lateral speeds of $0.4,0.5,0.6,0.7,0.8$ m/min are used corresponding to the first to the fifth columns of the confusion matrix labeled as $1$ to $5$. We observe that the accuracy of the lateral speed classification is consistently greater than 91% with a mean square error less than 1.5%.
@@
Case RMSE of $\hat{x}$ ACC1 of $\hat{v}$ ACC2 of $\hat{d}$ Fitness $g$
Case 1 (w. OPT Params) 0.0498$\pm$ 0.0042 97.37$\pm$0.7299.75$\pm$0.110.07849
Case 1 (w. RND Params) 0.0661$\pm$ 0.0120 93.47$\pm$7.3298.31$\pm$1.860.18397
Case 2 (w. OPT Params) 0.0712$\pm$ 0.0012 95.58$\pm$1.2199.37$\pm$0.210.12165
Case 2 (w. RND Params) 0.0645$\pm$ 0.0159 89.16$\pm$24.2398.89$\pm$1.90.1477
The satisfactory experimental results validate the proposed motion state estimation algorithm, which further indicates that the propeller wake features extracted from the ALL distributed pressure measurements are sufficiently rich for estimating the motion states of a leader propeller-driven underwater robot in a leader-follower formation.
IV-E Comparison Study
To demonstrate the effectiveness of our proposed estimation algorithm, we conducted a comparative experiment, comparing our approach with two traditional flow sensing methods from the literature including the frequency analysis-based method [27, 28] and the learning-based method [46, 47, 26]. Specifically, the frequency analysis-based method employs short-time Fourier transform (STFT) to extract localized time-varying frequency features from pressure measurement time series. These features are then inputted into a two-dimensional convolutional neural network to generate estimated states of relevant wake flows. The learning-based method adopts the one-dimensional convolutional neural network (1DCNN) to directly train upon the raw pressure measurement time series.
Figure 6 shows the experimental results of the statistical estimation accuracy regarding the three motion states of interest. The comparison reveals that our proposed flow estimation method consistently outperforms the conventional flow sensing algorithms across different estimation states and experimental conditions. Specifically, RMSE of the estimated lateral displacement $\hat{x}$ decreases by 75% and 30%, compared to the conventional frequency analysis-based and learning-based methods, respectively. Similarly, ACC of the estimated traveling speed $\hat{v}$ increases by 68% and 45%, and ACC of the estimated direction of motion $\hat{d}$ increases by 49% and 27%, compared to frequency analysis-based and learning-based methods, respectively. These results illustrate the significant advantage of the proposed method over conventional flow sensing approaches in addressing the unique propeller wake estimation challenge.
@@ V-A Length of Time Series on Estimation Performance
Figure 7 shows the experimental results of the influence of the time-series sequence length on the estimation performance represented by the corresponding evaluation metrics. We observe that as the sequence length increases, the RMSE of the lateral displacement decreases slightly indicating a gentle performance improvement. On the other hand, the classification accuracy of both the traveling speed and the motion of direction increases considerably when the sequence length increases. Overall, the longer length of the time-series is used in the estimation algorithm, the higher the motion estimation performance is achieved, which we conjecture comes from the extra flow dynamics information embedded in the added sampled data. However, longer sequence length leads to more time delay in the estimation process, thus possibly hindering real-time control tasks. Comprehensively considering the estimation performance and the real-time requirement of the ALL sensing system, the selection of the sequence length of the sensor measurement time series is expected to be resolved with respect to different applications. Particularly, $sl=64$ is adopted in this paper.
@@ V-B Optimized Task Weights on Estimation Performance
We conducted a comparative experiment using WOA-based optimized weights and randomly selected weights to assess their influence on overall estimation performance. In the comparative experiment, we trained 10 estimation models based on 10 sets of randomly generated task weights and another 10 estimation models based on the WOA optimized weights. We compared and analyzed the averaged fitness value $\bar{g}$ over 10 estimation models and the statistical state estimation accuracy between using random and optimized weights.
Table I illustrates the experimental results. We observe a significant decrease in fitness $g$ using optimized weights compared to using random weight, by roughly 57% and 17% in Case 1 and Case 2, respectively. Furthermore, in Case 1, we notice a significant enhancement in estimation performance with optimized weights. Specifically, there is an improvement in estimation accuracy of 4% in speed $v$, 1.4% in direction of motion $d$, and 24.6% in displacement $x$. However, the estimation accuracy for displacement $x$ in Case 2 decreases by 0.6%, which we conjecture is attributed to the highly-coupled nature of multi-objective optimization. In addition, the standard deviations of the estimation accuracy for all three states are reduced significantly by at least 65% with optimized weights, indicating a more robust and consistent estimation performance across all motion states and testing conditions. Therefore, optimizing the algorithm to obtain optimization weights is crucial.
V-C WOA vs GA and PSO
To justify our choice of the optimization algorithm, we conducted a comparative experiment, comparing WOA with classical Genetic Algorithm (GA) and Particle Swarm Optimization (PSO) algorithms. We adjusted some hyperparameters to expedite the process, e.g., reducing training steps to 100 from 200 and batch size to 512 from 64. For each optimization algorithm, we conducted five repeated tests varying only the initial weight values while keeping all other optimizer parameters unchanged. Fig. 8 shows the experimental results of averaged fitness value and computational time over the five repeated tests. While the fitness value of WOA is approximately 0.2% lower than that of GA and 0.03% higher than that of PSO, the WOA algorithm significantly reduces optimization time by approximately 7% and 8% over GA and PSO, respectively. Comprehensively considering the network complexity and extended computational time, we selected WOA for its effectiveness and efficiency as the task weight optimizer in this paper.
Figure 8: Experimental results of the averaged fitness value $\bar{g}$ and the computational time consumption, comparing WOA with GA and PSO in estimation model training. While WOA exhibits similar fitness value, it significantly reduces computational time, thus making WOA a preferred choice for optimizing network parameters in this paper.
@@ V-C WOA vs GA and PSO
V-D Uniqueness in Propeller Wake Sensing
Table II presents a comparison between our proposed method and others from the literature, focusing on the generation method of the flow field and whether multiple motion states are simultaneously estimated. Existing studies primarily address flow fields generated by periodic body undulation and tail flapping such as in bioinspired robotic fish, while our study deals with the dynamic and complex wake flow generated by high-speed rotating propellers. Moreover, our method enables simultaneous estimation of multiple motion states through a multi-output network design combining regression and classification operations. Therefore, we consider the problem addressed in this paper unique, significantly different from the existing studies.
V-E Limitations and Mitigation Measures
We would like to briefly discuss the limitations of the proposed propeller wake sensing method and explore mitigation measures to enhance its real-world application. First, the data-driven algorithm relies on pre-collected training data, so conducting more experiments is necessary to expand the dataset and cover more types of complex motions. Secondly, the horizontal sensor arrangement limits the sensing of motion states to the corresponding plane. Therefore, enhancing ALL with more distributed pressure sensors is critical for improving its spatial sensing capabilities. Thirdly, in this study, ALL was fixed on the experimental platform without considering sensor noise and pressure perturbation caused by its motion. A viable solution is to integrate inertia measurement units and fuse sensor measurements to mitigate such influences.
@@ VI Conclusions
This paper investigated the challenging problem of estimating the relative motion states of a propeller-driven leader underwater robot in the leader-follower formation. Making a hypothesis that the ALL pressure measurements provide partial but sufficiently rich information of the propeller wake of a leader underwater robot, we proposed a novel motion state estimation method that assimilated distributed pressure measurements sampled from the dynamic and complex propeller wake. Specifically, to estimate the lateral displacement, speed, and direction of motion of the leader propeller, a hybrid network that combined 1DCNN and BiLSTM was designed to extract the spatiotemporal characteristics of the flow field, followed by a multi-output network for state estimation. An experimental testbed was constructed. Extensive experimental tests were conducted, the results of which validated the proposed estimation method.